y a na the eflect on the equilibrium PHU IS Unce . 28) The demand curve is P = 700 - 20QD. The supply curve is P-300 + 200s. can some explain the steps for this
Answer #1

Ans: B ) 10;500

Explanation:

At equilibrium , QD = QS

700 - 20Q = 300 + 20Q

40Q = 400

Q = 400 / 40 = 10 ( Equilibrium quantity)

Now we put the value of Q in any equation to get the equilibrium price.

P = 700 - 20QD = 700 - ( 20 *10 ) = 700 - 200 = 500

P = 300 + 20QS = 300 + ( 20 *10 ) = 300 + 200 = 500